OpenStudy (anonymous):

y=2x^2 +8x write in vertex form

OpenStudy (anonymous):

Completing the square of 2x^2+8x will give you vertex form

OpenStudy (anonymous):

i tried that and i got a positive b when its negative how do we fix this?

OpenStudy (anonymous):

Show me how you completed the square

OpenStudy (anonymous):

y=(x+2)^2-8

OpenStudy (anonymous):

thats what i got

OpenStudy (anonymous):

That is correct.

OpenStudy (anonymous):

but is y= (x-2)^2 -8

OpenStudy (anonymous):

Nope the first one was right

OpenStudy (anonymous):

okay how about y=x^2-12x-3 how do we get this one?

OpenStudy (anonymous):

i got y=(x+6)^2 -39 when the answer is y=(x-6)^2 -39

OpenStudy (anonymous):

x^2-12x-3 (-12/2)^2=36 add and subtract 36 x^2-12x+36-36-3 factor first 3 terms (x-6)^2-39

OpenStudy (anonymous):

as a rule of thumb if the b value is negative then the h value is also going to be negative aswell right?

OpenStudy (anonymous):

What is b and what is h?

OpenStudy (anonymous):

-12 is b and h is -6

OpenStudy (anonymous):

Remember how you divide the middle term by 2 and then square it to get the number that you add and subtract? The number you get before squaring will be the number you use. -12/2=-6 and so it factored (x-6)^2

OpenStudy (anonymous):

so to get the h value we basically divide the b value by 2

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

alright thanks